Earleton Manufacturing Company has $3 billion in sales and $600,000,000 in fixed assets. Currently, the company's fixed assets are operating at 85% of capacity.

  1. What level of sales could Earleton have obtained if it had been operating at full capacity? Write out your answer completely. Round your answer to the nearest whole number.
    $  

  2. What is Earleton's target fixed assets/sales ratio? Round your answer to two decimal places.
    %

  3. If Earleton's sales increase 30%, how large of an increase in fixed assets will the company need to meet its target fixed assets/sales ratio? Write out your answer completely. Do not round intermediate calculations. Round your answer to the nearest whole number.
    $
Answer #1

a.

Sales at Full Capacity = 3,000,000,000/0.85

Sales at Full Capacity = $3,529,411,764.71

b.

Target FA/Sales = 600,000,000/3,529,411,764.71

Target FA/Sales = 0.17

c.

0.17 = FA/(1.30*3,000,000,000)

FA = $663,000,000

Change in FA = 663,000,000 - 600,000,000

Change in FA = $63,000,000